OpenStudy (anonymous):

determine the limit as x approaches zero. [1-cos(2x)]/[4x^2]

OpenStudy (anonymous):

\[(1-\cos(2x))\div(4x^2)\]

OpenStudy (anonymous):

hint-\[\cos2x=1-2\sin ^{2}x\] use this and get the answer.

OpenStudy (anonymous):

that is a part of what the answer says but how did you realise that cos2x equals this?

OpenStudy (anonymous):

is it an identity?

OpenStudy (anonymous):

oh okay thanks.. hmm

OpenStudy (anonymous):

you can also prove it.. cos(2x)=cos(x+x) use this and expand, you will land up as cos2x=1-2sin^2x=2cos^2x-1.

OpenStudy (anonymous):

oh okayy. So now i have -2sin^2x over 4x^2 im not sure what to do now

OpenStudy (anonymous):

cos2x=1-2sin^2x 2sin^2x=1-cos2x. now we have 2sin^2x over 4x^2

OpenStudy (anonymous):

lim x tending to 0, sinx/x=1 this is a result in limits did you know this before?

OpenStudy (anonymous):

hmm no I didnt.. im still a bit confused with the second last part. I thought that it would be 1-1-2sin^2x over 4x^2 ? i dont know why i went wrong. sorry

OpenStudy (anonymous):

okay.. cos2x=1-2sin^2x add 2sin^2x on both sides..what do you get?

OpenStudy (anonymous):

2sin^2x + cos2x = 1 ?

OpenStudy (anonymous):

good.. now to the equation which you've obtained, subtract cos2x from both sides. what do you get?

OpenStudy (anonymous):

2sin^2x = 1 - cos2x ohhh

OpenStudy (anonymous):

yes..now use the result lim x tending to zero, sinx/x=1 and get the answer.

OpenStudy (anonymous):

oh okay so it would be 2/4 or 1/2, but im confused why sinx/x = 1

OpenStudy (anonymous):

when x=0, what is sinx?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

yes..so when x=0, sinx/x=0/0 dividing a quantity sinx over x , at x=0 is undefined. do you agree?

OpenStudy (anonymous):

yep